1.    The lifetime of light bulbs produced by a company are normally distributed with mean 1000 hours and standard deviation 130 hours.

a)    The top 30% of all light bulbs should last at least how many hours?

b)    What is the probability that a randomly selected light bulb will last at least 900 hours?

c)    If we randomly select 8 light bulbs, what is the probability that the average of these light bulbs will last more than 1050 hours?


1
Expert's answer
2020-12-21T17:28:36-0500

Let "X=" the lifetime of light bulb in hours: "X\sim N(\mu, \sigma^2)"

Then "Z=\dfrac{X-\mu}{\sigma}\sim N(0, 1)"

Given "\mu=1000\ h, \sigma=130\ h"

a)

"P(X\geq x_1)=0.3"

"P(X\geq x_1)=1-P(X<x_1)"

"=1-P(Z<\dfrac{x_1-1000}{130})=0.3"

"P(Z<\dfrac{x_1-1000}{130})=0.7"

"\dfrac{x_1-1000}{130}\approx0.524401"

"x_1=1000+130(0.524401)=1068"

The top 30% of all light bulbs should last at least 1068 hours.


b)

"P(X\geq 900)=1-P(X<900)"

"=1-P(Z<\dfrac{900-1000}{130})"

"\approx1-P(Z<-0.76923077)"

"\approx1-0.220878=0.779122"

"0.779122"


c) Let "Y=" the average lifetime of these bulbs in hours: "Y\sim N(\mu, \sigma^2/n)"

Then "Z=\dfrac{Y-\mu}{\sigma/\sqrt{n}}\sim N(0, 1)"

Given "\mu=1000\ h, \sigma=130\ h, n=10"


"P(Y>1050)=1-P(Y\leq1050)"

"=1-P(Z\leq\dfrac{1050-1000}{130/\sqrt{10}})"

"\approx1-P(Z\leq1.21626064)"

"\approx1-0.888057=0.111943"

"0.111943"